# Paired Course and Dorm Allocation

### v3

@@ 3 Model
There is a continuum of students denoted by the interval $\mathcal{I}=[0,1]$ and endowed with the Lebesgue measure $\mu$. Let $i\in\mathcal{I}$ denote an individual student. There is a finite set $\mathcal{C}$ of courses and a finite set $\mathcal{D}$ of dorms (when thinking about other applications, $\mathcal{C}$ and $\mathcal{D}$ can be interpreted as arbitrary goods). Each $c\in\mathcal{C},d\in\mathcal{D}$ denotes an individual course or dorm, while $C\subset\mathcal{C},D\subset\mathcal{D}$ denotes a set of courses or dorms. Let $q:\mathcal{C}\cup\mathcal{D}\to\mathbb{R}$ denote a mapping from courses and dorms to how many students can enroll in a course or live in a dorm (so $q(c_{1})$ represents the mass of students that may be enrolled in course $c_{1}$). Student $i$ receives utility from a bundle of courses $C$ and a dorm $d$ according to
@@
so preferences are additively separable across courses and dorms. This assumption is meaningful beyond analytic tractability. At many institutions, course and dorm selection happen at different points in time: For example, undergraduates who are guaranteed housing need to be matched early on before graduate students can select their rooms whereas course enrollment cannot happen until professors figure out what and when they are teaching. As such, institutional constraints may restrict students from reporting richer preferences over courses and dorms. On the student’s side, preferences over one market may include expectations of the other market. One natural source of complementarities comes from students wanting to live close to their classes. In this case, even if students do not know which classes are available (both in general or for them to enroll in), they still know that large lectures will be held in large lecture halls and classes in certain departments will be near the department building. Thus, a statistics student may have “distance to the statistics building” as part of their utility from dorms without violating additive separability.
Each student has access to some finite and ordered set of signals $\mathcal{S}=\{1,...,S\}$ for some $S\in\mathbb{N}$. For example, simply asking students if they care more about courses or dorms is equivalent to setting $\mathcal{S}=\{1,2\}$; school administration may ask students to rank their relative preference between courses and dorms on a scale of one to ten, in which case $\mathcal{S}=\{1,2,...,10\}$. Let $\hat{s}(i)$ denote the signal that student $i$ reports. To account for ties, let $r^{c},r^{d}:\mathcal{I}\to[0,1]$ be two measurable and bijective functions (we will call them permutations of the unit interval). Let $R$ be a probability distribution over the space of all permutations of the unit interval from which $r^{c},r^{d}$ are drawn. [^5] Mathematically, it is impossible to draw uniformly from the space of all measurable, bijective functions from the unit interval to itself. As such, we can think of each agent having an improper prior over $r^{c},r^{d}$ while the true tie-breaking rule is chosen by the mechanism some other way but is unknown to the agents. Alternatively, given a sequence of finite markets using random tie-breaking over a finite number of permutations, we can recover our model via limit arguments developed in Carmona and Podczeck (2020). As regularity conditions, suppose:
@@
    Symmetric Tie-breaking: If $B$ is a set of permutations of the unit interval that is $R$-measurable, then for all $i,j\in\mathcal{S}$,
@@
    where
    $\sigma_{ij}(k)=\begin{cases}k&\text{ if }k\neq i,j\\ i&\text{ if }k=j\\ j&\text{ if }k=i\end{cases}$
    is the permutation of $i$ and $j$.
@@
    Run serial dictatorship for dorms ordering students via $-\hat{s}(i)+r^{d}(i)$
@@ 4 Welfare
As Proposition 4 holds for any paired serial dictatorship mechanism, a corollary is that even independent RSD [^8] Recall that independent RSD can be recovered from paired serial dictatorship if the signal space is a singleton. produces a result that is ex-ante envy-free. However, envy-freeness is a weak property that does not speak to overall efficiency. We will now discuss the efficiency of paired serial dictatorship.
4.1 Deterministic Efficiency
Students’ choices of signals are determined by their relative preferences between courses and dorms (Proposition 1) and how competitive their favorite courses and dorms are (Proposition 2). However, the final bundle a student receives is also mediated by tie-breaking, which could either harm or help coordination as seen in the motivating example. Without signaling, the mechanism can only resort to ad-hoc (random) tie-breaking which is uncorrelated with relative preferences. On the other hand where tie-breaking does not play a role and paired serial dictatorship induces a deterministic allocation (the courses and dorm each student is assigned to are constant in $r^{c},r^{d}$), the resulting outcome is Pareto efficient. In this case, allocating priorities via paired serial dictatorship is sufficient to resolve students’ conflicts of interest in the course and dorm assignment stage.
@@ 4.1 Deterministic Efficiency
Any equilibrium of paired serial dictatorship that induces a deterministic allocation is Pareto efficient.
The proof relies on showing that if some student were strictly better off under an alternative allocation that makes no students worse off, then that student must have had a profitable deviation from the original equilibrium. As such, even though we might expect externalities to arise from students switching priorities (one student reporting a higher signal to secure a spot in a certain class leads to some other student no longer being able to register for that class), each student best-responding to $\hat{s}$ in equilibrium fully mediates these externalities.
@@
so each student values $d$ more than $1+1/q(d)$ times more than their greatest difference in utility between any two bundles without $d$. Then, each student gets dorm $d$ with probability (at least) $1/q(d)$ if they report the lowest signal; that probability to get dorm $d$ outweighs all other considerations. As such, the unique Nash equilibrium (for any signal space) is for all students to report the lowest signal, leading back to full randomization. Hyper-competition for dorms is a practical concern that many (Stanford) students share. At worst, every student reporting the same signal leads back to independent randomization, while if a few students report something else, those students are better off under paired serial dictatorship than under independent RSD. This “free” improvement over independent RSD we get is a consequence of our large-market approach. In large, finite economies, this corresponds to the intuition that any one individual’s deviation has a vanishingly small probability of impacting any other student’s bundle.
@@ 5 Simulations
Next, we plot students’ equilibrium signal against the logarithm of their ratio of course preference to dorm preference, $\log(\lambda_{i}/\gamma_{i})$ as the logarithm of ratios is a symmetric distribution while ratio itself has a large right tail. Results are in Figure 3:
@@
In this particular setting, students exhibit bunching at high signals, suggesting that good courses are more competitive than good dorms. One hypothesis as to why this is the case is that course selection determines which four courses are chosen while dorm selection only determines a student’s one dorm. This force can be calibrated by changing the common value component of students’ utilities from courses or dorms. Students with higher log ratios tend to report higher signals, corroborating the intuition behind Proposition 1. However, Proposition 1 does not perfectly apply as students generally will not care more about the same courses or dorms than another student due to pseudo-random initialization of preferences.
Next, we consider changes in student welfare. As student utilities were initialized independent of one another, it is possible for one student to have twice as much ’utils’ from any bundle as another student, so comparisons of aggregate utility or even changes in utility are not meaningful in this setting. However, we can derive meaningful comparisons from analyzing each student’s percent change in welfare between the two mechanisms. Each individual student’s expected utility and standard deviation in utilities (across $r^{c},r^{d}$ draws) were generally improved; Out of $1,000$ students, $917$ saw an increase in expected utility while $985$ saw a decrease in standard deviation in utility. On average, student expected utility increased by $3.18\%$ while student standard deviations in utilities decreased by $57.28\%$. Histograms of percent changes in a student’s expected utility and standard deviations in utility from paired serial dictatorship versus independent RSD are plotted in Figure 4:
@@ 5.1 Expanding the Signal Space
What happens when students have access to a richer set of signals? In general, this can only help students better express preferences. To test this, we keep the same randomly initialized environment as before and only change the signal space to be $\{0,1,...,19\}$. Learning dynamics in this setting are similar to before: Total regret over time is comparable and a majority of students converge to basically a pure strategy: $970$ students play their most likely signal with probability greater than $0.9$, while $852$ students play their most likely signal with probability at least $1-10^{-9}$. Since students simply have twice as many signals as before, learning is slower in this setting. Equilibrium behavior still exhibits more competition for courses than dorms, but more separation is observed with students pooling between three signals (17, 18, 19) instead of the previous two (8, 9). Compared to before, more students also play lower signals outside of the signals that other students pool on. Note that Proposition 5 does not apply since the original equilibrium did not induce a deterministic allocation and exponential weight updating may not learn the same equilibria if there are multiple. Random sampling of $r^{c},r^{d}$ between trials and iterations may also lead to different equilibria being learned. Student behavior is plotted in Figure 5.
@@ 5.3 Bounded Rationality
where $4.5$ is halfway between the smallest signal 00 and the largest signal $9$ and $\log(\lambda_{i}/\gamma_{i})$ is symmetrically distributed around 00 with support approximately $[-6.5,6.5]$. Alternatively, we can think of this case as a heuristic test of Proposition 6 when the common value assumption fails. In the simulation setup, students do not have common values for courses and dorms, but under this specification of bounded rationality, students behave as if the common value assumption holds: Relative preferences are the only force driving students’ signals (formalized in Proposition 1) while the competitiveness of different courses or dorms (formalized in Proposition 2) is ignored,
@@ 6 Conclusion
  • $\psi^{c},\psi^{d}$ are (measurable) re-labelings of students given signals, $r^{c},r^{d}$ so student $i$ chooses courses before student $j$ if and only if $\psi^{c}(i)>\psi^{c}(j)$ and dorms before if and only if $\psi^{d}(i)>\psi^{d}(j)$.
@@
is still a finite union of a finite intersection of measurable sets, and is hence still measurable. In general, $c_{n}$ is the $n$th course to reach capacity, and $i_{n}$ is the first student for which $c_{n}$ (and by extension, $c_{1},...,c_{n-1}$) is unavailable.
@@
where $U_{a}(i,\hat{x})$ is the $a$-th component of $U(i,\hat{x})$ and the expectation is taken with respect to $r^{c},r^{d}\sim R$.
@@
[…]
@@
Step Three: Measurably
@@
be the set of player $i$’s best responses to $\hat{x}$, where $q\cdot U(i,\hat{x})=\sum_{a\in A}q(a)U_{a}(i,\hat{x})$ is the dot product between $q$ and $U(i,\hat{x})$.
@@
$e_{a}(a^{\prime})=\left\{\begin{matrix}1&\text{ if }a^{\prime}=a\\ 0&\text{ otherwise.}\\ \end{matrix}\right.$
@@
[…]
@@
$\Big{[}q^{n}\cdot U(i,\hat{x}^{n})-q^{0}\cdot U(i,\hat{x}^{0})\Big{]}\leq\|q^{n}-q^{0}\|_{L^{1}}\cdot\max_{a}\Big{[}U^{a}(i,\hat{x}^{n})-U^{a}(i,\hat{x}^{0})\Big{]}$
which converges to zero as $\max_{a}\Big{[}U^{a}(i,\hat{x}^{n})-U^{a}(i,\hat{x}^{0})\Big{]}$ is bounded (there are only a finite number of bundles any player can get) and $q^{n}\to q^{0}$ implies $\|q^{n}-q^{0}\|_{L^{1}}\to 0$.
@@
[…]
@@
where the expectation is taken with respect to $r^{c},r^{d}\sim R$. Suppose student $i$ cares more about the same courses than student $j$. For each fixed $r^{c},r^{d}$, let ${r^{c}}^{\prime},{r^{d}}^{\prime}$ be the tie-breaking rules that permute $i$ and $j$’s tie-break priority (so $r^{c}(i)={r^{c}}^{\prime}(j)$). By symmetry, ${r^{c}}^{\prime},{r^{d}}^{\prime}\sim R$ as well. Then, the courses students $i$ and $j$ receive are identical up to tie-breaking. As such,

### v4

@@ 3 Model
There is a continuum of students denoted by the interval $\mathcal{I}=[0,1]$ and endowed with the Lebesgue measure $\mu$ to model a large economy where each individual has zero aggregate effect. Let $i\in\mathcal{I}$ denote an individual student. There is a finite set $\mathcal{C}$ of courses and a finite set $\mathcal{D}$ of dorms (when thinking about other applications, $\mathcal{C}$ and $\mathcal{D}$ can be interpreted as arbitrary goods). Each $c\in\mathcal{C},d\in\mathcal{D}$ denotes an individual course or dorm, while $C\subset\mathcal{C},D\subset\mathcal{D}$ denotes a set of courses or dorms. Let $q:\mathcal{C}\cup\mathcal{D}\to\mathbb{R}$ denote a mapping from courses and dorms to how many students can enroll in a course or live in a dorm (so $q(c_{1})$ represents the mass of students that may be enrolled in course $c_{1}$). Student $i$ receives utility from a bundle of courses $C$ and a dorm $d$ according to
@@
so preferences are additively separable across courses and dorms. This assumption is meaningful beyond analytic tractability. At many institutions, course and dorm selection happen at different points in time: For example, undergraduates who are guaranteed housing need to be matched early on before graduate students can select their rooms, whereas course enrollment cannot happen until professors figure out what and when they are teaching. As such, institutional constraints may restrict students from reporting richer preferences over courses and dorms. On the student’s side, preferences over one market may include expectations of the other market. One natural source of complementarities comes from students wanting to live close to their classes. In this case, even if students do not know which classes are available (both in general or for them to enroll in), they still know that large lectures will be held in large lecture halls and classes in certain departments will be near the department building. Thus, a statistics student may have “distance to the statistics building” as part of their utility from dorms without violating additive separability.
Each student has access to some finite and ordered set of signals $\mathcal{S}=\{1,...,S\}$ for some $S\in\mathbb{N}$. For example, simply asking students if they care more about courses or dorms is equivalent to setting $\mathcal{S}=\{1,2\}$; school administration may ask students to rank their relative preference between courses and dorms on a scale of one to ten, in which case $\mathcal{S}=\{1,2,...,10\}$. Let $\hat{s}(i)$ denote the signal that student $i$ reports. To account for ties, let $r^{c},r^{d}:\mathcal{I}\to[0,1]$ be two measurable and bijective functions (we will call them permutations of the unit interval). Let $R$ be a probability distribution over the space of all permutations of the unit interval from which $r^{c},r^{d}$ are drawn. [^7] For a discussion on the existence of distributions over the set of all bijective functions from the unit interval from itself (such as the uniform distribution), see footnote 9 of Takahashi (2010). As regularity conditions, suppose:
@@
    Symmetric Tie-breaking: If $B$ is a set of permutations of the unit interval that is $R$-measurable, then for all $i,j\in\mathcal{I}$,
@@
    where $\sigma_{ij}$ is the permutation of $i$ and $j$.
@@
    Run serial dictatorship for dorms ordering students via $\hat{s}(i)+r^{d}(i)$
@@ 4 Welfare
As Proposition 4 holds for any paired serial dictatorship mechanism, a corollary is that even independent RSD [^10] Recall that independent RSD can be recovered from paired serial dictatorship if the signal space is a singleton. produces a result that is ex-ante envy-free. However, ex-ante envy-freeness is a weak property that does not speak to overall efficiency. It turns out that randomization, necessitated by tie-breaking, is the key barrier to establishing results about efficiency (and ex-post envy-freeness).
4.1 Deterministic Efficiency and Fairness
Students’ choices of signals are determined by their relative preferences between courses and dorms (Proposition 1) and how competitive their favorite courses and dorms are (Proposition 2). However, the final bundle a student receives is also mediated by tie-breaking, which could either harm or help coordination as seen in the motivating example. Without signaling, the mechanism can only resort to ad-hoc (random) tie-breaking which is uncorrelated with relative preferences. On the other hand, when tie-breaking does not play a role and paired serial dictatorship induces a deterministic allocation (the courses and dorm each student is assigned to are constant in $r^{c},r^{d}$), the resulting outcome is Pareto efficient. In this case, allocating priorities via paired serial dictatorship is sufficient to resolve students’ conflicts of interest in the course and dorm assignment stage.
@@ 4.1 Deterministic Efficiency and Fairness
Any equilibrium of paired serial dictatorship that induces a deterministic allocation is Pareto efficient and envy-free.
The proof relies on showing that if some student were strictly better off under an alternative allocation that makes no students worse off, then that student must have had a profitable deviation from the original equilibrium. As such, even though we might expect externalities to arise from students switching priorities (one student reporting a higher signal to secure a spot in a certain class leads to some other student no longer being able to register for that class), each student best-responding to $\hat{s}$ in equilibrium fully mediates these externalities. At deterministic allocations, the ex-ante distribution over allocations is degenerate, and Proposition 4 directly implies that deterministic allocations are envy-free.
@@
so each student values $d$ more than $1+1/q(d)$ times more than their greatest difference in utility between any two bundles without $d$. Then, each student gets dorm $d$ with probability (at least) $q(d)$ if they report the lowest signal; that probability to get dorm $d$ outweighs all other considerations. As such, the unique Nash equilibrium (for any signal space) is for all students to report the lowest signal, leading back to full randomization. Hyper-competition for dorms is a practical concern that many students share. At worst, every student reporting the same signal leads back to independent randomization, while if a few students report something else, those students are better off under paired serial dictatorship than under independent RSD. This “free” improvement over independent RSD we get is a consequence of our large-market approach. In large, finite economies, this corresponds to the intuition that any one individual’s deviation has a vanishingly small probability of impacting any other student’s bundle.
@@ 5 Simulations
Next, we plot students’ equilibrium signal against the logarithm of their ratio of course preference to dorm preference, $\log(\lambda_{i}/\gamma_{i})$ as the logarithm of ratios is a symmetric distribution but the distribution of ratios itself has a large right tail. Results are in Figure 3:
@@
In this particular setting, students exhibit bunching at high signals, suggesting that good courses are more competitive than good dorms. One hypothesis as to why this is the case is that course selection determines which four courses are chosen while dorm selection only determines a student’s one dorm. This force can be calibrated by changing the common value component of students’ utilities from courses or dorms. Students with higher log ratios tend to report higher signals, corroborating the intuition behind Proposition 1. However, Proposition 1 does not perfectly apply, as students generally will not care more about the same courses or dorms than another student due to pseudo-random initialization of preferences.
Next, we consider changes in student welfare. As student utilities were initialized independent of one another, it is possible for one student to have twice as much ’utils’ from any bundle as another student, so comparisons of aggregate utility or even changes in utility are not meaningful in this setting. However, we can derive meaningful comparisons from analyzing each student’s percent change in welfare between the two mechanisms. Each individual student’s expected utility and standard deviation in utilities (across $r^{c},r^{d}$ draws) were generally improved. Out of $1,000$ students, $917$ saw an increase in expected utility while $985$ saw a decrease in standard deviation in utility. On average, student expected utility increased by $3.18\%$ while student standard deviations in utilities decreased by $57.28\%$. Histograms of percent changes in a student’s expected utility and standard deviations in utility from paired serial dictatorship versus independent RSD are plotted in Figure 4:
@@ 5.1 Expanding the Signal Space
What happens when students have access to a richer set of signals? In general, this can only help students better express preferences. To test this, we keep the same randomly initialized environment as before and only change the signal space to be $\{0,1,...,19\}$. Learning dynamics in this setting are similar to before: Total regret over time is comparable and a majority of students converge to basically a pure strategy: $970$ students play their most likely signal with probability greater than $0.9$, while $852$ students play their most likely signal with probability at least $1-10^{-9}$. Since students simply have twice as many signals as before, learning is slower in this setting. Equilibrium behavior still exhibits more competition for courses than dorms, but more separation is observed with students pooling between three signals (17, 18, 19) instead of the previous two (8, 9). Compared to before, more students also play lower signals outside of the signals that other students pool on. Note that Proposition 5 does not apply since the original equilibrium did not induce a deterministic allocation and exponential weight updating may not learn the same equilibria if there are multiple. Random sampling of $r^{c},r^{d}$ between trials and iterations may also lead to different equilibria being learned. Student behavior is plotted in Figure 5:
@@ 5.3 Bounded Rationality
where $4.5$ is halfway between the smallest signal 0 and the largest signal $9$ and $\log(\lambda_{i}/\gamma_{i})$ is symmetrically distributed around 0 with support approximately [6.5,6.5][-6.5,6.5]. Alternatively, we can think of this case as a heuristic test of Proposition 6 when the common value assumption fails. In the simulation setup, students do not have common values for courses and dorms, but under this specification of bounded rationality, students behave as if the common value assumption holds: Relative preferences are the only force driving students’ signals (formalized in Proposition 1) while the competitiveness of different courses or dorms (formalized in Proposition 2) is ignored.
@@ 6 Conclusion
  • $\psi^{c},\psi^{d}$ are (measurable, bijective) re-labelings of students given signals, $r^{c},r^{d}$ so student $i$ chooses courses before student $j$ if and only if $\psi^{c}(i)>\psi^{c}(j)$ and dorms before if and only if $\psi^{d}(i)>\psi^{d}(j)$.
@@
$i_{c_{n+1}}=\sup\left\{i:\min_{c\in\mathcal{C}\setminus\{c_{1},...,c_{n}\}}\left(q(c)-e_{i_{n}}(c)-\nu\Big(\{i_{c_{n}}\leq j<i:c\in C_{j}(\mathcal{C}\setminus\{c_{1},...,c_{n}\})\Big)\right)>0\right\}$
@@
is still a finite union of a finite intersection of measurable sets, and is hence still measurable. In general, $c_{n}$ is the $n$th course to reach capacity, and $i_{c_{n}}$ is the first student for which $c_{n}$ (and by extension, $c_{1},...,c_{n-1}$) is unavailable.
@@
[…]
@@
Let $\hat{x}_{a}$ denote the $a$-th component of $\hat{x}$. The auxiliary utility function $U$ depends only on $\int_{\mathcal{I}}\hat{x}_{a}(i)d\mu(i)$ for $a\in A$.
@@
Step Three: Measurability
@@
[…]
@@
[…]
@@
$\int_{\mathcal{I}}\hat{x}^{n}_{a}(i)d\mu(i)=\int_{\mathcal{I}}\hat{x}^{n}(i)\cdot e_{a}d\mu(i)\to\int_{\mathcal{I}}\hat{x}^{0}(i)\cdot e_{a}d\mu(i)=\int_{\mathcal{I}}\hat{x}^{0}_{a}(i)d\mu(i)$
@@
$\Big[q^{n}\cdot U(i,\hat{x}^{n})-q^{0}\cdot U(i,\hat{x}^{0})\Big]\leq\|q^{n}-q^{0}\|_{L^{1}}\cdot\max_{a}\left[\left|U_{a}(i,\hat{x}^{n})-U_{a}(i,\hat{x}^{0})\right|\right]$
which converges to zero as $\max_{a}\Big[\left|U_{a}(i,\hat{x}^{n})-U_{a}(i,\hat{x}^{0})\right|\Big]$ is bounded (there are only a finite number of bundles any player can get) and $q^{n}\to q^{0}$ implies $\|q^{n}-q^{0}\|_{L^{1}}\to 0$.
@@
It must be that $i$ chooses before $t_{\hat{x}^{0}}$ (but after $t_{\hat{x}^{n}}$). Then, there is an open interval of time between $i$ choosing and $t_{\hat{x}^{0}}$. As $t_{\hat{x}}$ varies continuously in $\left(\int_{\mathcal{I}}\hat{x}_{a}(i)d\mu(i)\right)_{a\in A}$, $\hat{x}^{n}\to\hat{x}^{0}$ implies that $t_{\hat{x}^{n}}\to t_{\hat{x}^{0}}$. As such, the bundle $i$ receives under $\hat{x}^{n}$ will eventually be no worse (and potentially better) than the bundle they receive under $\hat{x}^{0}$. Since this holds for any $r^{c},r^{d}$, as $n\to\infty$ the probability (with respect to nature drawing $r^{c},r^{d}$) of $i$ receiving a better bundle under $\hat{x}^{0}$ than $\hat{x}^{n}$ tends to zero. As utilities are bounded, this in turn implies that $q^{*}\cdot\Big[U(i,\hat{x}^{0})-U(i,\hat{x}^{n})\Big]$ is non-positive for sufficiently large $n$.
@@
where the expectation is taken with respect to $r^{c},r^{d}\sim R$. Suppose student $i$ cares more about the same courses than student $j$. For each fixed $r^{c},r^{d}$, let ${r^{c}}^{\prime},{r^{d}}^{\prime}$ be the tie-breaking rules that permute $i$ and $j$’s tie-break priority (so $r^{c}(i)={r^{c}}^{\prime}(j)$). By symmetry, ${r^{c}}^{\prime},{r^{d}}^{\prime}\sim R$ as well. Then, the courses students $i$ and $j$ receive are identical up to tie-breaking as they have the same ordering over bundles of courses. As such,